OpenStudy (anonymous):

At which point will the graphs of the equations 2x + y = 8 and x - y = 4 intersect?

OpenStudy (anonymous):

Set them both equal to Y, and then set the functions equal

OpenStudy (anonymous):

y = -2x = 8 and y = x-4

OpenStudy (anonymous):

\[-2x - 8=x-4\]

OpenStudy (johnweldon1993):

I'm assuming you meant y = -2x + 8 and y = x - 4

OpenStudy (anonymous):

thats what i wrote

OpenStudy (anonymous):

yeah thats what i meangt

OpenStudy (luigi0210):

-2x+8=x-4 solve

OpenStudy (johnweldon1993):

As Luigi is showing...you then set them equal to each other and solve for 'x' -2x + 8 = x - 4

OpenStudy (anonymous):

x = 4

OpenStudy (johnweldon1993):

right, now you can take that 'x' value...and plug it into either of your equations y = -2x + 8 y = x - 4 replace the 'x' with (4) and see what 'y' equals...

OpenStudy (anonymous):

y = 0

OpenStudy (anonymous):

ohhh i get it thanks

OpenStudy (johnweldon1993):

Right, so these graphs will intersect at (4,0) No problem